OpenStudy (anonymous):

is there away to figure out what f(x) is if we know the expression for f(1/x)? I'm just gonna make up a random one. Say f(1/x) = (x^2 + 1)/x. What's f(x)?

myininaya (myininaya):

\[f(\frac{1}{x})=\frac{x^2+1}{x} \\ \] replace 1/x with u or replace x with 1/u

myininaya (myininaya):

\[f(u)=\frac{(\frac{1}{u})^2+1}{\frac{1}{u}}\]

myininaya (myininaya):

now f(x)=?

OpenStudy (primeralph):